OpenStudy (anonymous):

Find the range of 5/8 , 3/4 , 1 1/8, 1 1/4

OpenStudy (anonymous):

the range of this data set is all x between (smallest number) and (highest number)

ganeshie8 (ganeshie8):

lets make the denominators common.. so we can compare them

ganeshie8 (ganeshie8):

\[\frac{5}{8}, \frac{3}{4}, \frac{9}{8}, \frac{5}{4} \] is this the dataset given after converting mixed fractions ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

\[\frac{5}{8}, \frac{6}{8}, \frac{9}{8}, \frac{10}{8}, \]

ganeshie8 (ganeshie8):

since their denominators are same. we can compare them now.. lowest is with the lowest numerator, highest is with the highest numerator. range = [highest - lowest ] can you compute ?